# Tuan Duc Ngo UMass Amherst &Ashkan Mirzaei Snap Inc. &Guocheng Qian

### v1

@@ 5 Conclusion
In this work, we presented an efficient framework for dense 3D video tracking that significantly improves the runtime of DELTA ngo2024delta while preserving its strong performance. We proposed a coarse-to-fine tracking algorithm that progressively increases spatial coverage across iterations, combined with a learnable interpolation module for dense supervision and a faster 4D correlation computation. These improvements yield 5100×5-100$\times$5 - 100 × speedup over previous approaches, making our approach more suitable for real-time applications. Nonetheless, our method inherits common limitations of data-driven tracking: it is trained on synthetic data and may struggle under fast motion, severe occlusions, or poor depth estimation, which affect 2D/3D tracking accuracy in complex real-world scenes.
@@ Appendix A More Results
We further evaluate our method on the TAP-Vid3D benchmark using depth maps predicted by ZoeDepth [4]. Results are summarized in Tab. 7.
@@
TAP-Net [10] 38.5 54.4 80.6 33.0 48.6 78.8 54.6 68.3 87.7
MFT [43] 39.6 60.4 72.7 47.3 66.8 77.8 - - -
PIPs [18] 31.7 53.7 72.9 42.2 64.8 77.7 15.7 28.4 77.1
OmniMotion [64] - - - 46.4 62.7 85.3 69.5 82.5 90.3
TAPIR [12] 49.6 64.2 85.0 56.2 70.0 86.5 54.2 69.8 84.4
CoTracker2 [32] 48.7 64.3 86.5 60.6 75.4 89.3 63.1 77.0 87.8
DOT [34] 48.4 63.8 85.2 60.1 74.5 89.0 77.1 87.7 93.3
BootsTAPIR [11] 54.6 68.4 86.5 61.4 73.6 88.7 70.8 83.0 89.9
TAPTR [37] 49.0 64.4 85.2 63.0 76.1 91.1 - - -
TAPTRv2 [36] 49.7 64.2 85.7 63.5 75.9 91.4 - - -
LocoTrack [8] 52.9 66.8 85.3 63.0 75.3 87.2 69.0 83.2 89.5
SpatialTracker [69] 50.1 65.9 86.9 61.1 76.3 89.5 63.5 77.6 88.2
SceneTracker [62] - 66.5 - - 71.8 - - 73.3 -
@@
Figure 6: Qualitative comparison of dense 3D tracking between SceneTracker [62], SpaTracker [69], DELTA [44], and our method. We track every pixel from the first frame through 3D space, with moving objects highlighted in rainbow colors. Our approach accurately captures foreground motion, preserves stable backgrounds, and operates significantly faster with a 5×5$\times$5 × to 100×100$\times$100 × speedup compared to prior methods.
@@ A.2 Dense 3D Tracking
Figure 6 shows visual comparisons of dense 3D tracking between our method and baselines. Our model produces highly stable and accurate trajectories, clearly outperforming SceneTracker [62] and SpaTracker [69], and achieving results on par with the strong DELTA [44] baseline.
@@ B.1 Further Analysis of Strategies for Accelerating Tracking
1) Spatial downsampling of RGB input: applied with spatial reduction factors of 2$\times$, 4$\times$, and 8$\times$.
2) Temporal subsampling of input frames: applied with temporal reduction factors of 2$\times$, 4$\times$, and 8$\times$.
3) Trajectory point subsampling: applied with trajectory subsampling ratios of 4$\times$, 16$\times$, and 64$\times$.
@@ B.2 Spectral Analysis of the optical flow predictions
To analyze the spatial frequency characteristics of optical flow, we follow prior work [54] on spectral analysis using the two-dimensional discrete cosine transform (2D DCT) [1]. Given a flow field $\mathcal{F}_{flow}\in\mathbb{R}^{H\times W\times 2}$, we process the horizontal and vertical components separately by dividing each into non-overlapping blocks of size $B\times B$. We then apply the type-II 2D DCT to each block, which projects the flow values onto a set of cosine basis functions at different spatial frequencies. Within each block, we arrange the coefficients into a one-dimensional sequence using the standard JPEG zigzag ordering [61], which traverses from low to high spatial frequencies. To summarize the overall spectral content of the flow, we average these zigzag-ordered coefficients across all blocks and across both flow components (horizontal and vertical). This produces a compact frequency profile that captures the distribution of motion energy across spatial scales.
@@
Table 9: Comparison of cost reduction strategies on the Kubric3D val set [44]. All methods use 1 iteration unless noted; full-resolution outputs are obtained using either bilinear or nearest-neighbor interpolation.
@@
Strategy 2×2$\times$2 × 4×4$\times$4 × 8×8$\times$8 ×

### v2

@@ 5 Conclusion
In this work, we presented an efficient framework for dense 3D video tracking that significantly improves the runtime of DELTA ngo2024delta while preserving its strong performance. We proposed a coarse-to-fine tracking algorithm that progressively increases spatial coverage across iterations, combined with a learnable interpolation module for dense supervision and a faster 4D correlation computation. These improvements yield $5-100\times$ speedup over previous approaches, making our approach more suitable for real-time applications. Nonetheless, our method inherits common limitations of data-driven tracking: it is trained on synthetic data and may struggle under fast motion, severe occlusions, or poor depth estimation, which affect 2D/3D tracking accuracy in complex real-world scenes.
Acknowledgements
Evangelos Kalogerakis has received funding from the European Research Council (ERC) under the Horizon research and innovation programme (Grant agreement No. 101124742).
@@ Appendix A More Results
We further evaluate our method on the TAP-Vid3D benchmark using depth maps predicted by ZoeDepth [zoedepth]. Results are summarized in Tab. 7.
@@
TAP-Net [tapvid] 38.5 54.4 80.6 33.0 48.6 78.8 54.6 68.3 87.7
MFT [mft] 39.6 60.4 72.7 47.3 66.8 77.8 - - -
PIPs [pips] 31.7 53.7 72.9 42.2 64.8 77.7 15.7 28.4 77.1
OmniMotion [omnimotion] - - - 46.4 62.7 85.3 69.5 82.5 90.3
TAPIR [tapir] 49.6 64.2 85.0 56.2 70.0 86.5 54.2 69.8 84.4
CoTracker2 [cotracker] 48.7 64.3 86.5 60.6 75.4 89.3 63.1 77.0 87.8
DOT [dot] 48.4 63.8 85.2 60.1 74.5 89.0 77.1 87.7 93.3
BootsTAPIR [bootstap] 54.6 68.4 86.5 61.4 73.6 88.7 70.8 83.0 89.9
TAPTR [taptr] 49.0 64.4 85.2 63.0 76.1 91.1 - - -
TAPTRv2 [taptrv2] 49.7 64.2 85.7 63.5 75.9 91.4 - - -
LocoTrack [cho2024local] 52.9 66.8 85.3 63.0 75.3 87.2 69.0 83.2 89.5
SpatialTracker [spatialtracker] 50.1 65.9 86.9 61.1 76.3 89.5 63.5 77.6 88.2
SceneTracker [scenetracker] - 66.5 - - 71.8 - - 73.3 -
@@
Figure 6: Qualitative comparison of dense 3D tracking between SceneTracker [scenetracker], SpaTracker [spatialtracker], DELTA [ngo2024delta], and our method. We track every pixel from the first frame through 3D space, with moving objects highlighted in rainbow colors. Our approach accurately captures foreground motion, preserves stable backgrounds, and operates significantly faster with a $5\times$ to $100\times$ speedup compared to prior methods.
@@ A.2 Dense 3D Tracking
Figure 6 shows visual comparisons of dense 3D tracking between our method and baselines. Our model produces highly stable and accurate trajectories, clearly outperforming SceneTracker [scenetracker] and SpaTracker [spatialtracker], and achieving results on par with the strong DELTA [ngo2024delta] baseline.
@@ B.1 Further Analysis of Strategies for Accelerating Tracking
[…]
@@ B.2 Spectral Analysis of the optical flow predictions
To analyze the spatial frequency characteristics of optical flow, we follow prior work [skorokhodov2025improving] on spectral analysis using the two-dimensional discrete cosine transform (2D DCT) [ahmed2006discrete]. Given a flow field $\mathcal{F}_{flow}\in\mathbb{R}^{H\times W\times 2}$, we process the horizontal and vertical components separately by dividing each into non-overlapping blocks of size $B\times B$. We then apply the type-II 2D DCT to each block, which projects the flow values onto a set of cosine basis functions at different spatial frequencies. Within each block, we arrange the coefficients into a one-dimensional sequence using the standard JPEG zigzag ordering [wallace1991jpeg], which traverses from low to high spatial frequencies. To summarize the overall spectral content of the flow, we average these zigzag-ordered coefficients across all blocks and across both flow components (horizontal and vertical). This produces a compact frequency profile that captures the distribution of motion energy across spatial scales.
@@
Table 9: Comparison of cost reduction strategies on the Kubric3D val set [ngo2024delta]. All methods use 1 iteration unless noted; full-resolution outputs are obtained using either bilinear or nearest-neighbor interpolation.
@@
Strategy $2\times$ $4\times$ $8\times$